# Online Residual Learning from Offline Experts for Pedestrian Tracking

### v1

@@ V-C Simulation Results
We present results for two individual pedestrians on two different scenes, namely the ‘Hyang0’ and the ‘Hyang3’, as they are termed in the SDD. More specifically, we present instantaneous and cumulative error plots, as well as predicted trajectory plots on each scene. In these plots, the prediction resulting from our proposed method is compared to the methods listed in Table I, namely the purely online prediction, where we perform RLS prediction directly on the target $r_{t}$, without considering any offline experts, the prediction resulting from considering only the offline predictions as experts, without online adaptation, as well as the best offline one which is defined as the one that has the smallest Average Displacement Error (ADE): $\mathrm{ADE}_{i}=\frac{1}{T}\sum_{t=1}^{T}\ell_{t}(\hat{r}_{t,i})$.

### v2

@@ V-C Simulation Results
We present results for two individual pedestrians on two different scenes, namely the ‘Hyang0’ and the ‘Hyang3’, as they are termed in the SDD. More specifically, we present instantaneous and cumulative error plots, as well as predicted trajectory plots on each scene. In these plots, the prediction resulting from our proposed method is compared to the following methods listed in Table I: i) the online method, where we perform RLS prediction directly on the target $r_{t}$ without considering any offline experts, ii) the offline experts method, where we consider directly the raw offline predictions as experts, and we combine them online using Algorithm 2, as well as iii) the best offline one which returns the prediction with the smallest Average Displacement Error (ADE): $\mathrm{ADE}_{i}=\frac{1}{T}\sum_{t=1}^{T}\ell_{t}(\hat{r}_{t,i})$.